If 50% of 2y is 12, what is y^2?
A)36
B) 6
C)16
D)144

Answer by ikleyn(52781)

.

If 50% of 2y is 12, then 2y is 24.


Hence,  y is  24/2 = 12;  then  y^2 = 12^2 = 144.     ANSWER